Title of manuscript: Force line breaks with \\

A. A. Author and B. B. Author Lines break automatically or can be forced with \\

Authors’ institution and/or address This line break forced with \\

C. C. Author

Second author institution and/or address This line break forced with \\ (Received 1 January 1989)

The author will not know the received date when the compuscript is first submitted; production will insert this. Every article includes an abstract. The abstract is a concise summary of the work covered at length in the main body of the article. It is used for secondary publications and for information retrieval purposes. Valid PACS numbers should be entered after the abstract is finished, using the \pacs command. Valid PACS appear here. \pacs{} should always be input, even if empty.

I. FIRST-LEVEL HEADING: THE LINE BREAK WAS FORCED VIA \\

Here is the first sentence in Sec. I, demonstrating section cross-referencing. Note that this sample file was run with the eqsecnum option selected.

A. Second-level heading: The line break was forced via \\

Here is the first sentence in Sec. I A, demonstrating section cross-referencing. The command \narrowtext will make the text this width. The command \widetext will make the text the width of the full page, as on page 3. A blank input line tells TEX that a new paragraph begins.

Reference citations in text use the command \cite. In the reference section of this paper each reference is “tagged” by a string (in curly brackets). The proper form for citing in text is \cite{string}, and the result is shown here [1, 2]. We will cite other people [1, 3] and journals here. We also cite other people again (Refs. [1] and [3]). It is worth mentioning that REVTEX “collapses” lists of reference numbers where possible. We now cite everyone together [1–3], and once again (Refs. [1–3]).

II. DISPLAYED EQUATIONS

A. Another second-level heading

1. Third-level heading: The line break was forced via \\

Here is the first sentence in Sec. II A 1, demonstrating section cross-referencing. In LaTEX there are many different ways to display equations, and a few preferred ways are noted below.

a. Fourth-level heading: Single-line equations. Below we have indented, single-line equations with numbers; this is the most common type of equation in Physical Review:

χ_{+}(p) < ∼ [2|p|(|p| + p_{z})]^{−1/2} (\begin{matrix}|p| + p_{z} \\ px + ip_{y} \end{matrix} ) , (2.1)

\{11234567890abc123αβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} , (2.2)

Note the open one in Eq. (2.2). If the equation is a little wider, the equation number automatically moves down to the next line:

{\{abc1234567890abc1234αβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} . \atop (2.3)}

When the \label command is used [cf. input for Eq. (2.2)], the equation can be referred to in text without your knowing the equation number that TEX will assign to it.

Math will indent 18.4 pt by default; not all equations will fit within a narrow column this way. Sometimes an equation will fit on one line when it is set flush-left with the margin. \FL is the command to set the math flush left:

{\{ab12345678abc123456abcdefαβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} , \atop (2.4)}

{ϵ_{μ}^{∗}(p) → c(V)D_{V} ∑_{τ} c_{τ}^{f} ū(f)P_{τ} γ_{μ}v(f̄) [ϵ_{j}l_{i}ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}) . \atop (2.5)}

Use \FL before every equation that you do not want indented. If you do not use \FL, equations will automatically indent.

The command \FR is also available, and sets equations flush right.

If you have a single-line equation that you don’t want numbered, you can use the \[, \] format:

g^{+}g^{+} → g^{+}g^{+}g^{+}g^{+} …, q^{+}q^{+} → q^{+}g^{+}g^{+} ….

2. Multiline equations

Multiline equations are obtained by using the

\begin{eqnarray}, \end{eqnarray} format. Use the

\nonumber command at the end of each line where you do not want a number:

M = ig_{Z}^{2}(4E_{1}E_{2})^{1/2}(l_{i}^{2})^{−1}δ_{σ_{1},−σ_{2}}(g_{σ_{2}}^{e})^{2}χ_{−σ_{2}}(p_{2})
×[ϵ_{j}l_{i}ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}),  (2.6)
∑ |M_{g}^{viol}|^{2} = g_{S}^{2n−4}(Q^{2}) N^{n−2}(N^{2} − 1)

× (∑_{i<j}) ∑_{perm} \frac{1}{S_{12}} \frac{1}{S_{12}} ∑_{τ} c_{τ}^{f} . (2.7)

If you wish to set a multiline equation without any line numbers, you can use the \begin{eqnarray*},

\end{eqnarray*} format:

∑ |M_{g}^{viol}|^{2} = g_{S}^{2n−4}(Q^{2}) N^{n−2}(N^{2} − 1)

× (∑_{i<j}) ( ∑_{perm} \frac{1}{S_{12}S_{23}S_{n1}} ) \frac{1}{S_{12}} .

To obtain numbers not normally produced by the automatic numbering, use the \eqnum{#1} command, where #1 is the desired equation number. For example, to get an equation number of (2.7′),

{g^{+}g^{+} → g^{+}g^{+}g^{+}g^{+} …, q^{+}q^{+} → q^{+}g^{+}g^{+} …. \atop (2.7′)}

A few notes on \eqnum. The \eqnum must come before the \label, if any. The numbering set with \eqnum is transparent to the automatic numbering in REVTEX; therefore, you must know the number ahead of time, and must make sure that the number set with \eqnum stays in step with the automatic numbering. \eqnum works with both single-line and multiline equations. You could, if you wished, do all the numbering in a paper manually

with \eqnum.

Enclosing single-line and multiline equations in

\begin{mathletters} and \end{mathletters} will

produce a set of equations that are “numbered” with letters, as shown in Eqs. (2.8a) and (2.8b) below:

\{abc123456abcdefαβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} , (2.8a)

M = ig_{Z}^{2}(4E_{1}E_{2})^{1/2}(l_{i}^{2})^{−1}(g_{σ_{2}}^{e})^{2}χ_{−σ_{2}}(p_{2})
×[ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}).  (2.8b)

3. Wide equations

The equation that follows is set in a wide format, i.e., it spans across the full page. The wide format is reserved for long equations that cannot be easily broken into four lines or less:

R^{(d)} = g^{3} cos θ_{W} δ_{σ_{1},−σ_{2}} [g_{σ_{2}}^{e} (\frac{[Γ^{Z}(3, 21)]_{σ_{1}}}{Q_{12}^{2} − M_{W}^{2}} + \frac{[Γ^{Z}(13, 2)]_{σ_{1}}}{Q_{13}^{2} − M_{W}^{2}} ) + x_{W} Q_{e} (\frac{[Γ^{γ}(3, 21)]_{σ_{1}}}{Q_{12}^{2} − M_{W}^{2}} + \frac{[Γ^{γ}(13, 2)]_{σ_{1}}}{Q_{13}^{2} − M_{W}^{2}} )] . (2.9)

This is typed so you can see that the output is in wide format. (Since there is no input line between \end{equation} and this paragraph, there is no paragraph indent for this paragraph.) We have

R^{(f)} = −g^{3}δ_{σ_{1},σ_{2}}M_{Z}^{2}x_{W}^{2} (\frac{g_{σ_{2}}^{e}D_{Z}}{cos θ_{W}} − Q_{e}D_{γ} cos θ_{W} ) (\frac{[ϵ_{3}]_{σ_{1}}}{Q_{12}^{2} − M_{W}^{2} /ξ} ϵ_{1} · ϵ_{2} + \frac{[ϵ_{2}]_{σ_{1}}}{Q_{13}^{2} − M_{W}^{2} /ξ} ϵ_{1} · ϵ_{3}) . (2.10)

III. CROSS-REFERENCING

REVTEX will automatically number sections, equations, figure captions, and tables. In order to reference them in text, use the \label and \ref commands.

The \label command appears following a section heading or within a equation, figure caption, or table; the \ref command appears in text where citation is to occur. We will refer to the first figure (Fig. 1) here. We can refer to the “late figure” also (Fig. 2).

References to figures: Fig. 1, Fig. 2, Fig. 3, and Fig. 4. References to tables: Table I, Table II, Table III, Table

IV, Table V, Table VI; Physical Review style requires that the initial citation of figures or tables be in numerical order in text, so don’t cite Fig. 3 until you’ve cited Fig. 2. See Style and Notation Guide.

ACKNOWLEDGMENTS

We wish to acknowledge the tireless efforts of our typesetters...

If a section does not have a number (like the Acknowledgments section), use the command \nonum. For the Acknowledgments section you can also use the command \acknowledgments to produce the heading.

APPENDIX A: APPENDIXES

To start appendixes, you should use either the

\appendix{#1} or the \unletteredappendix{#1} com-

mand. Use the latter if there is only one appendix in a paper, and the former if there is more than one. #1 represents the heading of the appendix, and can be left empty.

APPENDIX B: A LITTLE MORE ON APPENDIXES

Note the equation number in an appendix:

E = mc^{2}, (B1)

1. A subsection in an appendix

You can use a subsection or subsubsection in an appendix. Note the numbering: we are now in Appendix B 1.

Note the equation numbers in this appendix, produced with the mathletters environment:

E = mc, (B2a)

E = mc^{2}, (B2b)

E > ∼ mc^{3}. (B2c)

They turn out to be Eqs. (B2a), (B2b), and (B2c).

∗ Also at Physics Department, XYZ University.

[1] A. Smith and B. Doe, J. Chem. Phys. 76, 4056 (1982).

[2] C. Jones, J. Chem. Phys. 68, 5298 (1978).

[3] C. Jones and A. Smith, J. Chem. Phys. 72, 3416 (1980); 73, 5168 (1980); 72, 4009 (1980).

FIG. 1. A figure caption. The figure captions are automatically numbered.

FIG. 2. The “late figure.” This figure was inserted when the paper was finished. Since the figures are automatically numbered, no renumbering in text was necessary. All that needed to be done was to type the caption in the proper place and cite the figure in text.

FIG. 3. A figure caption. Figures will be reduced to an appropriate size by the production staff upon receipt.

FIG. 4. A figure caption. The labels you give tables and figures can be descriptive (as that of Fig. 1, which has a \label{autonum}) or they can reflect their numerical order, as that of this figure (\label{fig4}).

TABLE I. This is a narrow table, which occupies the width of a narrow column. The table captions are automatically numbered.

One  Two  Three
one  two  three
one  two  three

TABLE II. A medium table. Pseudopotential core radii fit to the observed bulk modulus. Also given are the radius r_{0} of the atomic volume and the ratio κr_{0} of r_{0} to the screening length.

r_{c} (Å)  r_{0} (Å)  κr_{0}  r_{c} (Å)  r_{0} (Å)  κr_{0}
Cu  0.800  1.410  2.550  Sn  0.680  1.870  3.700
Ag  0.990  1.590  2.710  Pb  0.450  1.930  3.760
Au  1.150  1.590  2.710  Ca  0.750  2.170  3.560
Mg  0.490  1.760  3.200  Sr  0.900  2.370  3.720
Zn  0.300  1.520  2.970  Li  0.380  1.730  2.830
Cd  0.530  1.710  3.160  Na  0.760  2.110  3.120
Hg  0.550  1.780  3.220  K  1.120  2.620  3.480
Al  0.230  1.580  3.240  Rb  1.330  2.800  3.590
Ga  0.310  1.670  3.330  Cs  1.420  3.030  3.740
In  0.460  1.840  3.500  Ba  0.960  2.460  3.780
Tl  0.480  1.890  3.550

TABLE III. A wide table. Two alternative occupations of special positions by KMnCL_{3} ions in the two space groups D_{4h}^{1} and D_{4h}^{1}. For a special value of the x and y parameters, a set of special positions may split into two sets of special positions of higher symmetry.

D_{4h}^{1}  D_{4h}^{5}
Ion  1st alternative  2nd alternative  lst alternative  2nd alternative
K  (2e) + (2f)  (4i)^{a}  (2c) + (2d)  (4f)^{a}
Mn  (2g)^{a}  (a) + (b) + (c) + (d)  (4e)^{a}  (2a) + (2b)
Cl  (a) + (b) + (c) + (d)  (2g)^{a}+(2h)^{a}  (a) + (2b)  (4e)^{a}
He  (8r)^{a}  (4j)^{a}  (4g)^{a}
Ag  (4k)^{a}  (4h)^{a}

aThe z parameter of these positions is z ∼ \frac{1}{4} .

TABLE IV. Another wide table. This shows decimal and implied-decimal alignments. Predecimal zeros required according to style.

One  Two  Three  Four  Five
one  two  three  four  five
He  2  2.77234  45672  0.69
C^{a}  C^{b}  12537.64  37.66345  86.37

aSome tables require footnotes. bSome tables need more than one footnote.

TABLE V. A “late table.” This table was added after most of the paper had been completed. Since the tables are automatically numbered, no renumbering in text was necessary. This table was added to show the use of the @ specifier for lining things up.

Align by .  Multiple alignments
23.890 12  23.890 12± 0.002
12 323 123 223 ±344
0.834 390 12  80.80 ± 45.3416

aHere we cite Ref. [2].

TABLE VI. The Poisson ratio defined as the ratio of lateral contraction to longitudinal expansion for uniaxial stress. Experimental values are given for comparison.

σ  σ
Predicted  Observed^{a}  Predicted  Observed^{a}
Cu  0.48  0.36  Al  0.47  0.33
Ag  0.48  0.37  Tl  0.47  0.35
Au  0.48  0.36  Sn  0.46  0.33
Mg  0.47  0.35  Pb  0.46  0.40–0.45
Zn  0.47  0.25  Pb  0.49  0.43
K  0.49  0.44

aHere we also cite a reference (Ref. [2]).